OpenStudy (anonymous):

Buckstars sells two types of coffee: one that costs $5.00 per pound and another that costs $6.20 per pound. How many pounds of the $5.00 per pound coffee should be mixed in with the $6.20 per pound coffee in order to have 30 pounds of a blend worth $5.50?

OpenStudy (blacksteel):

Suppose we use x pounds of the $5.00 coffee and y pounds of the $6.20 coffee. Then we set up a system of 2 equations. We know we want a total of 30 lb, so: x+y = 30 We also know we want the resulting coffee to cost $5.50 per pound, so the proportion of $5.00 to $6.20 coffee should allow for this. The coffee will be made up of x/30 of the $5.00 coffee and y/30 of the $6.20 coffee: x/30 * 5.00 + y/30 * 6.20 = 5.50 Solve the system: x+y = 30 x/30 * 5 + y/30 * 6.2 = 5.5 x = 30 - y 5x + 6.2y = 165 5(30-y) + 6.2y = 165 150 - 5y + 6.2y = 165 1.2y + 150 = 165 1.2y = 15 y = 12.5 x+y = 30 x + 12.5 = 30 x = 17.5 We can check this by plugging the numbers into the equations: x+y = 30 17.5 + 12.5 = 30 GOOD x/30*5 + y/30*6.2 = 5.5 17.5/30 * 5 + 12.5/30 * 6.2 = 5.5 2.9166666 + 2.58333333 = 5.5 GOOD